aaron2020:

Emilio has to run 71 2 miles every week to train for the football team. If he runs 11 2 miles every day, how many days will it take him to run 71 2 miles? It will take Emilio days to run 71 2 miles.

Extrinix:

Can you reword this question please?

AZ:

\( 7 \dfrac{1}{2}\) and \( 1 \dfrac{1}{2}\) right?

AZ:

If so, all you need to do is divide First change \( 7 \dfrac{1}{2}\) into an improper fraction like this 

AZ:

Can you do the same thing for \( 1 \dfrac{1}{2}\) ?

AZ:

And then so every day he's running 1 1/2 miles and how many days will it take to run 7 1/2 miles so all you have to do is divide you need to use the improper fractions that you got earlier so 7 1/2 would be 15/2 and what do you get for 1 1/2 ? so you have to divide 15/2 divided by whatever 1 1/2 is as an improper fraction and this is the formula for dividing fractions \( \dfrac{a}{b} \div \dfrac{c}{d} ~~~= ~~~\dfrac{a}{b} \times \dfrac{d}{c} ~~~= ~~~\dfrac{a\times d}{b \times c}\)